OpenStudy (anonymous):

what is the area of a rectangle with the length of 6 2/5 and a width of 4 1/6 inches A. 26 2/3 B. 24 1/15 C. 10 17/30 D. 13 1/3

OpenStudy (anonymous):

A=lw A=6 2/5 X 4 1/6 A=80/3 A=26 2/3